# Coupled Cluster Downfolding Theory in Simulations of Chemical Systems on Quantum Hardware

### v1

@@ III.5 Solvers
Inspired by ADAPT-VQE’s adaptive selection and the generator-coordinate method (GCM), ADAPT-GCIM uses the same procedure to iteratively choose UCC excitation generators from the same operator pool as fermionic ADAPT-VQE. Rather than optimizing their parameters, it employs these fixed generators to construct a non-orthogonal, overcomplete many-body basis. Projecting the system Hamiltonian into this basis yields an effective Hamiltonian whose generalized eigenvalue problem produces ground- and excited-state energiesZheng et al. (2023, 2024). Although ADAPT-GCIM requires no parameter optimization, a few classical optimization rounds can greatly accelerate convergenceZheng et al. (2024). We denote this variant ADAPT-GCIM$(x,y)$, where up to $y$ optimization rounds occur for every $x$ ansatz-selection iterations.
@@ III.6 Experiments on quantum computers
To confirm the feasibility of the downfolded Hamiltonians and qubit-ADAPT-VQE implementation on real quantum devices, we performed one-shot evaluations of qubit-ADAPT-VQE circuits on a trapped-ion (Quantinuum H1-1) and two superconducting (IBM $ibm\_marrakesh$, $ibm\_kingston$) devices for benzene (cc-pVTZ) and FBP, as previously listed in Table 3. Additionally, Appendix D details device specifications. Due to credit limits, we did not run ZNE experiments on H1‑1 for FBP.
@@
In experiments on Quantinuum emulator and hardware, ZNE method, originally developed in Ref. 109, greatly improved the solution quality. Specifically, our implementation of ZNE is modifyied from the “MP after” method in Ref. 110 for trapped-ion hardware. We inserted extra 2-qubit gates only in the second half of each circuit. The noise factor equals the ratio of 2-qubit gates in the amplified versus unamplified ansatz. Figure 4 shows linear regressions of energy versus noise factor; surprisingly, the extrapolated energy from the H1‑1 device aligns more closely with the noiseless value than that from the emulator. Each circuit ran with 1,024 shots, for a total of 119,808 shots for benzene (including noise-amplified circuits in ZNE) and 60,416 shots for FBP (unamplified only, since ZNE was not performed).
@@ IV Outlook
In this work, we utilized a simple form of the effective Hamiltonian based on rank-1 and rank-2 (and the simplest terms contributing to rank-3) commutator expansions, using external singly and doubly excited cluster amplitudes stemming from the conventional CCSD simulations. We restricted the many-body components to one- and two-body interactions. This simplified approach leaves room for future improvements in energy accuracy. For example, in the recent study Singh et al. (2025), we demonstrated that the inclusion of three-body interactions in effective Hamiltonians leads to significant improvements in ground-state properties, especially in strongly correlated regimes.
@@ VIII Data and code availability
The scalar, one- and two-body elements of the effective Hamiltonians were obtained using the parallel DUCC implementation developed in the ExaChem electronic structure code (https://github.com/ExaChem/exachem). The many-body form of the DUCC(3) effective Hamiltonian was derived and implemented in ExaChem using the symbolic algebra tool SymGen (https://github.com/npbauman/SymGen). The effective Hamiltonians of the molecular systems are hosted on a public repository (https://github.com/npbauman/DUCC-Hamiltonian-Library.git).
The data and the code for the related numerical simulation are available in the GitHub repository (https://github.com/pnnl/QuGCM).
@@ Appendix B CC downfolding: many-body structure of the effective Hamiltonian
The CC downfolding theory describes a quantum system composed of two or more interacting components. In the simplest case, where the quantum system $S$ consists of two interacting parts, $X$ and $Y$, represented by active and inactive spin-orbitals, respectively. CC downfolding can be used to construct an effective Hamiltonian for subsystem $X$ that reproduces the energy of the entire system $S$.
@@
where $n_{\rm act}$ is active-electrons number operator, the diagrammatic analysis (as shown in Fig. 5) reveals that CC downfolding facilitates explicit correlation between the active space and its orthogonal complement. The diagram in Fig. 5 illustrates a scenario where the number of active-space electrons is preserved; however, in the intermediate states, electrons can hop between the active and inactive spin-orbitals.
Figure 5: (a) Typical Hugenholtz diagram contributing to the one-body part of the effective Hamiltonian. Gray ovals represent amplitudes defining the $\sigma_{\rm ext}$/$\sigma_{\rm ext}^{\dagger}$ operators. The black dot represents antisymmetrized vertices corresponding to pairwise interactions in the bare Hamiltonian. The letter indices $I,J,K$ / $A,B$ ($\bar{k}$ / $\bar{c},\bar{d}$) denote occupied/virtual active (inactive) spin-orbitals. This diagram contributes to the $a_{B}^{\dagger}a_{A}$ excitation, which commutes with the active-electrons number operator, $n_{\rm act}=\sum_{P}a_{P}^{\dagger}a_{P}$, where $P$ is a generic active spin-orbital index. The intermediate states $S_{1}$ and $S_{2}$, however, involve configurations in the Hilbert space that correspond to the transfer of electrons between active (embedded system) and inactive (environment) spin-orbitals needed to correlate interacting subsystems $X$ and $Y$ defined by active and inactive orbitals. The diagram is represented in action of singly excited determinant $|\Phi_{M}^{A}\rangle=a_{A}^{\dagger}a_{M}|\Phi\rangle$. (b) a symbolic representation of triple excitation, with respect to the reference function $|\Phi\rangle$, defining the intermediate $S_{1}$ state. One can notice that the active electron in occupied active spin-orbital $J$ is promoted to virtual inactive spin-orbital $\bar{c}$, (c) a symbolic representation of triple excitation, with respect to the reference function $|\Phi\rangle$, defining the intermediate $S_{2}$ state.
@@ Appendix C UCCGSD ansatz and VQE and GCM algorithms
$\Gamma=\sum_{p,q}\gamma_{p}^{q}a_{p}^{\dagger}a_{q}+\sum_{p<q,r<s}\gamma_{pq}^{rs}a_{p}^{\dagger}a_{q}^{\dagger}a_{s}a_{r}.$ (9)
@@
In this work, we adopted the qubit-ADAPT-VQE algorithmTang et al. (2021b) to further simplify the ansatz of the variational algorithm. Specifically, instead of using the linear combination of Pauli strings mapped from fermionic excitation operators, the operator pool in qubit-ADAPT-VQE only selects the individual Pauli strings from the summations on exponents in Eqs. (12) and (13) with an odd number of Pauli $Y$. The long $Z$ tails of the Pauli strings are also removed for hardware efficiency. This greatly reduces the circuit depth of each Pauli rotation and the number of 2-qubit gates. In addition, Ref. 115 proves that using a minimum operator pool leads to a number of ansatz that grows linearly with the number of qubits, which reduces the total number of measurements for each ADAPT iteration.
@@
While highly successful in its traditional applications, the full GCM formulation can become computationally demanding, especially when dealing with a large number of generator coordinates or complex basis sets. This computational bottleneck has motivated the development of more flexible and computationally efficient alternatives. One such development is the Generator-Coordinate-Inspired Method (GCIM). GCIM retains the spirit of GCM by constructing a variational wave function as a linear combination of non-orthogonal states, but it offers greater flexibility in the choice and construction of these “basis states” (or configurations) compared to the strictly collective-coordinate-dependent generator states of traditional GCM. This allows for a more general exploration of the many-body Hilbert space, potentially incorporating configurations that are not easily mapped to simple collective coordinates, while still leveraging the variational principle.
@@
As in Ref. 52, ADAPT-GCIM can employ the same operator pool, which consists of excitation operators in Eq. (10) and Eq. (11), and the state preparation is the same as in Eq. (14). Rather than optimizing parameters for a VQE objective function, it fixes all parameters a pre-defined constant, then iteratively expands a non-orthogonal, overcomplete basis set using prepared states {|$\{\ket{\psi_{HF}},\ket{\psi^{(1)}},\dots\}$. Projecting the Hamiltonian $H$ onto this basis gives an effective Hamiltonian $\mathbf{H}$ and overlap matrix $\mathbf{S}$, leading to the generalized eigenvalue problem $\mathbf{H}f=\epsilon\mathbf{S}f$, whose smallest eigenvalue can be an accurate estimation of the ground-state energy of $H$, while some of the other eigenvalues correspond to the excitation-state energies.
@@ Appendix D Hardware Specifications
Table 5 lists the specifications of the quantum hardware we used, as reported by the vendors. Quantinuum’s data are averages obtained via benchmarking tests in Ref. 116. IBM’s data comes from its daily calibration, and the values in the table are averages that exclude failed gates (error rate equals 1) ibm (2025).
@@
A brute-force extension of QPE and VQE algorithms towards larger number of qubits
On the other hand lot of algorithmic developmnets in quantum information science focus on chemistry as a “killer app”. Quantum approaches QPE - changes the way how quantum chemistry simulations will be perfromed - paving the road towards the hypotsesis driven.
Integrate the most appealing feature of two computational approaches and use
The problem of practical utilization of quantum computing in simulations of chemical systems and processes has sparked an intensive effort in the last decade towards efficient design of wave function ansatzes and optimization algorithms that can enable effectively mapping of idealized algorithms into qubit registers.
Convergence of various computational models. Integrating HPC and cloud computing resources with quickly evolving. This is suppose to address several challanges in practical quantum simulations of chemical systems and processes
Quantum computing can be currently used in the context of small-size active spaces, which in optimistic case can only capture static correlation effects.
challenges to be solved: inability of incorporating in a systematic way dynamical correlations effects that are beyond the small-active space
Dense Hamiltonains are used!!!!
Realistic basis sets are used!!!!
Recent advances in CC pave the way for practical utilizatoin of quantum architectures in perfroming chemistry simulations. The hybrid
Assuring extensivity at the early stages of the quantum algorithms development for chemistry , which paves the way towards further extension of these methodologies to the extensive systems, where not assuring size-extensivity/size-consisteny may have an detrimental effect on the accuracy of the simulations.
quantum centric or hybrid
@articlepople1976theoretical, title=Theoretical models incorporating electron correlation, author=Pople, John A and Binkley, J Stephen and Seeger, Rolf, journal=International Journal of Quantum Chemistry, volume=10, number=S10, pages=1–19, year=1976, publisher=Wiley Online Library
@articlepople1999quantum, title=Quantum chemical models (Nobel lecture), author=Pople, John A, journal=Angewandte Chemie International Edition, volume=38, number=13-14, pages=1894–1902, year=1999, publisher=Wiley Online Library
@articlepople1999nobel, title=Nobel lecture: Quantum chemical models, author=Pople, John A, journal=Reviews of Modern Physics, volume=71, number=5, pages=1267, year=1999, publisher=APS

### v3

@@ III.5 Solvers
Inspired by ADAPT-VQE’s adaptive selection and the generator-coordinate method, ADAPT-GCIM uses the same procedure to iteratively choose UCC excitation generators from the same operator pool as fermionic ADAPT-VQE. Rather than optimizing their parameters, it employs these fixed generators to construct a non-orthogonal, overcomplete many-body basis. Projecting the system Hamiltonian into this basis yields an effective Hamiltonian whose generalized eigenvalue problem produces ground- and excited-state energiesZheng et al. (2023, 2024). Although ADAPT-GCIM requires no parameter optimization, a few classical optimization rounds can greatly accelerate convergenceZheng et al. (2024). We denote this variant ADAPT-GCIM$(x,y)$, where up to $y$ optimization rounds occur for every $x$ ansatz-selection iterations.
@@ III.6 Experiments on quantum computers
To confirm the feasibility of the downfolded Hamiltonians and qubit-ADAPT-VQE implementation on real quantum devices, we performed one-shot evaluations of qubit-ADAPT-VQE circuits on a trapped-ion (Quantinuum H1-1) and two superconducting (IBM $ibm\_marrakesh$, $ibm\_kingston$) devices for benzene (cc-pVTZ) and FBP (cc-pVDZ), as previously listed in Table 3. Additionally, Appendix D details device specifications. Due to credit limits, we did not run ZNE experiments on H1‑1 for FBP.
@@
In experiments on Quantinuum emulator and hardware, ZNE method, originally developed in Ref. 126, greatly improved the solution quality. Specifically, our implementation of ZNE is modifyied from the “MP after” method in Ref. 127 for trapped-ion hardware. We inserted extra 2-qubit gates only in the second half of each circuit. The noise factor equals the ratio of 2-qubit gates in the amplified versus unamplified ansatz. Figure 4 shows linear regressions of energy versus noise factor; surprisingly, the extrapolated energy from the H1‑1 device aligns more closely with the noiseless value than that from the emulator. Each circuit ran with 1,024 shots, for a total of 119,808 shots for benzene (including noise-amplified circuits in ZNE) and 60,416 shots for FBP (unamplified only, since ZNE was not performed).
@@ IV Outlook
In this work, we utilized a simple form of the effective Hamiltonian based on rank-1 and rank-2 (and the simplest terms contributing to rank-3) commutator expansions, using external singly and doubly excited cluster amplitudes stemming from the conventional CCSD simulations. We restricted the many-body components to one- and two-body interactions. This simplified approach leaves room for future improvements in energy accuracy. For example, in the recent study Singh et al. (2025), we demonstrated that the inclusion of three-body interactions in effective Hamiltonians leads to significant improvements in ground-state properties, especially in strongly correlated regimes.  Additionally, problems defined by downfolded Hamiltonians can be effectively integrated with other quantum solvers recently discussed in the literature Weaving et al. (2025); Zhao et al. (2025).
@@ VIII Data and code availability
The scalar, one- and two-body elements of the effective Hamiltonians were obtained using the parallel DUCC implementation developed in the ExaChem electronic structure code (https://github.com/ExaChem/exachem). The many-body form of the DUCC(3) effective Hamiltonian was derived and implemented in ExaChem using the symbolic algebra tool SymGen (https://github.com/npbauman/SymGen). The effective Hamiltonians of the molecular systems, as well as other data and the code for related numerical simulations, are hosted on a public repository (https://github.com/npbauman/DUCC-Hamiltonian-Library).
@@ Appendix B CC downfolding: many-body structure of the effective Hamiltonian
The CC downfolding theory describes a quantum system composed of two or more interacting components. In the simplest case, the quantum system $S$ consists of two interacting parts, $X$ and $Y$, represented by active and inactive spin-orbitals, respectively. CC downfolding can be used to construct an effective Hamiltonian for subsystem $X$ that reproduces the energy of the entire system $S$.
@@
where $n_{\rm act}$ is active-electrons number operator, the diagrammatic analysis (as shown in Fig. 5) reveals that CC downfolding facilitates explicit correlation between the active space and its orthogonal complement. The diagram in Fig. 5 illustrates a scenario where the number of active-space electrons is preserved in the initial and final states; however, in the intermediate states, electrons can hop between the active and inactive spin-orbitals.
Figure 5: (a) Typical Hugenholtz diagram contributing to the one-body part of the effective Hamiltonian. Gray ovals represent amplitudes defining the $\sigma_{\rm ext}$/$\sigma_{\rm ext}^{\dagger}$ operators. The black dot represents antisymmetrized vertices corresponding to pairwise interactions in the bare Hamiltonian. The letter indices $I,J,K$ / $A,B$ ($\bar{k}$ / $\bar{c},\bar{d}$) denote occupied/virtual active (inactive) spin-orbitals. This diagram contributes to the $a_{B}^{\dagger}a_{A}$ excitation, which commutes with the active-electrons number operator, $n_{\rm act}=\sum_{P}a_{P}^{\dagger}a_{P}$, where $P$ is a generic active spin-orbital index. The intermediate states $S_{1}$ and $S_{2}$, however, involve configurations in the Hilbert space that correspond to the transfer of electrons between active (embedded system) and inactive (environment) spin-orbitals needed to correlate interacting subsystems $X$ and $Y$. The diagram is represented in action of singly excited determinant $|\Phi_{K}^{A}\rangle=a_{A}^{\dagger}a_{K}|\Phi\rangle$. (b) a symbolic representation of triple excitation, with respect to the reference function $|\Phi\rangle$, defining the intermediate $S_{1}$ state. One can notice that the active electron in occupied active spin-orbital $J$ is promoted to virtual inactive spin-orbital $\bar{c}$, (c) a symbolic representation of triple excitation, with respect to the reference function $|\Phi\rangle$, defining the intermediate $S_{2}$ state.
@@ Appendix C UCCGSD ansatz and VQE and GCM algorithms
$\Gamma=\sum_{p,q}\gamma_{p}^{q}a_{p}^{\dagger}a_{q}+\sum_{p<q,r<s}\gamma_{pq}^{rs}a_{p}^{\dagger}a_{q}^{\dagger}a_{s}a_{r},$ (9)
@@
In this work, we adopted the qubit-ADAPT-VQE algorithmTang et al. (2021b) to further simplify the ansatz of the variational algorithm. Specifically, instead of using the linear combination of Pauli strings mapped from fermionic excitation operators, the operator pool in qubit-ADAPT-VQE only selects the individual Pauli strings from the summations on exponents in Eqs. (12) and (13) with an odd number of Pauli $Y$. The long $Z$ tails of the Pauli strings are also removed for hardware efficiency. This greatly reduces the circuit depth of each Pauli rotation and the number of 2-qubit gates. In addition, Ref. 133 proves that using a minimum operator pool leads to a number of ansatz that grows linearly with the number of qubits, which reduces the total number of measurements for each ADAPT iteration.
@@
While highly successful in its traditional applications, the full GCM formulation can become computationally demanding, especially when dealing with a large number of generator coordinates or complex basis sets. This computational bottleneck has motivated the development of more flexible and computationally efficient alternatives. One such development is the Generator-Coordinate-Inspired Method. GCIM retains the spirit of GCM by constructing a variational wave function as a linear combination of non-orthogonal states, but it offers greater flexibility in the choice and construction of these “basis states” (or configurations) compared to the strictly collective-coordinate-dependent generator states of traditional GCM. This allows for a more general exploration of the many-body Hilbert space, potentially incorporating configurations that are not easily mapped to simple collective coordinates, while still leveraging the variational principle.
@@
As in Ref. 69, ADAPT-GCIM can employ the same operator pool, which consists of excitation operators in Eq. (10) and Eq. (11), and the state preparation is the same as in Eq. (14). Rather than optimizing parameters for a VQE objective function, it fixes all parameters a pre-defined constant, then iteratively expands a non-orthogonal, overcomplete basis set using prepared states {|$\{\ket{\psi_{HF}},\ket{\psi^{(1)}},\dots\}$. Projecting the Hamiltonian $H$ onto this basis gives an effective Hamiltonian $\mathbf{H}$ and overlap matrix $\mathbf{S}$, leading to the generalized eigenvalue problem $\mathbf{H}f=\epsilon\mathbf{S}f$, whose smallest eigenvalue can be an accurate estimation of the ground-state energy of $H$, while some of the other eigenvalues correspond to the excitation-state energies.
@@ Appendix D Hardware Specifications
Table 5 lists the specifications of the quantum hardware we used, as reported by the vendors. Quantinuum’s data are averages obtained via benchmarking tests in Ref. 134. IBM’s data comes from its daily calibration, and the values in the table are averages that exclude failed gates (whose error rates equal to 1) ibm (2025).